OpenStudy (anonymous):

6x + 8 = –4 * Solve for x A.) -12 B.) 12 C.) -2 D.) 2

OpenStudy (alexandervonhumboldt2):

6x+8=-4 6x+8-8=-4-8 6x=-12 divide both sides by 6: 6x/6=-12/6 x=-2

OpenStudy (anonymous):

can u help with 8 more

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

–5x + 8 = –32 * Solve for x A.) -40 B.) 40 C.) -8 D.) 8

OpenStudy (anonymous):

nvm got it